A student records the repair cost for 51 randomly selected stereos. A sample mean of $57.03 and standard deviation of $16.74 are subsequently computed. Determine the 99% confidence interval for the mean repair cost for the stereos. Assume the population is normally distributed. Step 1 of 2 : Find the critical value that should be used in constructing the confidence interval. Round your answer to three decimal places.
Answer #1

Df = n - 1 = 51 - 1 = 50

From T table,

t critical value at 0.01 significance level with 50 df = 2.678